### v1

# Table-Text Alignment: Explaining Claim Verification Against Tables in Scientific Papers

Abstract
Scientific claim verification against tables typically requires predicting whether a claim is supported or refuted given a table. However, we argue that predicting the final label alone is insufficient: it reveals little about the model’s reasoning and offers limited interpretability. To address this, we reframe table–text alignment as an explanation task, requiring models to identify the table cells essential for claim verification. We build a new dataset by extending the SciTab benchmark with human-annotated cell-level rationales. Annotators verify the claim label and highlight the minimal set of cells needed to support their decision. After the annotation process, we utilize the collected information and propose a taxonomy for handling ambiguous cases. Our experiments show that (i) incorporating table alignment information improves claim verification performance, and (ii) most LLMs, while often predicting correct labels, fail to recover human-aligned rationales, suggesting that their predictions do not stem from faithful reasoning. [^1] Our data and code are available at https://github.com/Alab-NII/SciTabAlign
Table-Text Alignment: Explaining Claim Verification Against Tables in Scientific Papers
Xanh Ho,1 Sunisth Kumar,2 Yun-Ang Wu,3∗ Florian Boudin,4 Atsuhiro Takasu,1  and Akiko Aizawa1,2 1National Institute of Informatics, Japan           2The University of Tokyo, Japan 3National Taiwan University           4JFLI, CNRS, Nantes Université, France {xanh, takasu, aizawa}@nii.ac.jp           sunisth@g.ecc.u-tokyo.ac.jp r11944072@csie.ntu.edu.tw           florian.boudin@univ-nantes.fr
**footnotetext: Research conducted during internship at NII, Japan.
1 Introduction
Claim verification against tables requires models to determine whether a natural language claim is supported or refuted based on structured tabular data. Several benchmarks have been proposed in the general domain, such as TabFact Chen et al. (2020), INFOTABS Gupta et al. (2020), and FEVEROUS Aly et al. (2021), primarily focusing on Wikipedia tables. However, tables in scientific papers pose additional challenges: they are often denser, more structured, and require domain-specific reasoning.
Two datasets have recently addressed this task in the scientific domain: SEM-TAB-FACTS (SEM; Wang et al., 2021), which includes both claim verification and cell-level evidence selection, and SciTab Lu et al. (2023), which focuses solely on claim verification. While SEM includes an alignment component, its claims are crowd-generated and simplified, limiting their representativeness. SciTab, in contrast, uses naturally occurring claims but lacks explicit annotations that explain why a given label is correct.
We argue that label prediction alone, as in SciTab, is not enough. It fails to reveal whether a model truly understands the table content, nor does it provide interpretable reasoning. For both evaluation and practical use, models need to go beyond classification and provide explanations grounded in tabular evidence.
From the perspective of scientific reading tools Lo et al. (2023), table–text alignment is also crucial. It allows readers to quickly locate which parts of a table are referenced in the text, improving comprehension and accelerating the reading process. Such alignments could directly support scientific workflows by making tabular evidence more accessible and interpretable.
To address these limitations, we reframe table–text alignment as an explanation task for scientific claim verification. Specifically, we extend the SciTab dataset with human-annotated cell-level rationales. For each claim–table pair, annotators verify the claim label and highlight the minimal set of table cells needed to support the decision.
During annotation, we frequently encountered ambiguous cases in claim interpretation and evidence selection. To capture these edge cases systematically, we introduce a taxonomy of five ambiguity types in scientific table-based verification: (i) Table Conversion Errors, (ii) Additional Context Requirements, (iii) Unexpected Claim Types, (iv) Subjective Adjectives, and (v) Unclear Claims.
We use our dataset to evaluate various types of large language models (LLMs), including table-based models, open-source LLMs, and closed-source LLMs. Our experiments also incorporate three different prompting strategies. On average, our human-annotated cell-level rationales help improve the performance of the claim label prediction task. The results show that while models achieve high macro-F1 scores on the claim label prediction task, their performance on the cell selection task remains low—even for advanced models like GPT-4o. The highest score, 50.8, is achieved by Qwen 2.5 72B using CoT prompting. Further analysis of the correlation between the two tasks reveals that although LLMs often correctly predict the claim label, their ability to identify the corresponding explanation cells is still limited.
2 Related Work
Claim verification has been studied across multiple domains, including news Wang (2017), Wikipedia Thorne et al. (2018); Jiang et al. (2020), scientific literature Wadden et al. (2020); Ou et al. (2025), and medicine Kotonya and Toni (2020); Vladika et al. (2024). Beyond plain text, recent efforts have extended claim verification to structured or multimodal evidence, including tables Chen et al. (2020); Lu et al. (2023), figures Akhtar et al. (2024), knowledge graphs Kim et al. (2023) and multimodal data Yang et al. (2025b).
Among table-based datasets, SEM Wang et al. (2021) and TabEvidence Gupta et al. (2022) are most related to our work. However, SEM features simplified, crowd-generated claims, while TabEvidence is limited to two-column Wikipedia tables, lacking the complexity of scientific tables. Recent frameworks like Chain-of-Table Wang et al. (2024) and Dater Ye et al. (2023) include evidence selection steps, but only report label accuracy, without evaluating the relevance or quality of the selected evidence, limiting trust in their predictions.
In contrast, our work emphasizes explanation via alignment, explicitly evaluating whether the model selects the correct table cells needed for verification, providing a more faithful and interpretable assessment of reasoning.
3 Dataset Creation
3.1 Base Dataset: SciTab
We build on SciTab Lu et al. (2023), the only available dataset for claim verification against scientific tables with naturally occurring claims. SciTab is derived from SciGen Moosavi et al. (2021), a table-to-text generation dataset in which each sample consists of a scientific table and its corresponding textual description.
The dataset contains 1,224 claim–table pairs: 457 supported, 411 refuted, and 356 not enough information (NEI). Supported claims are sourced from original paper content, while refuted and NEI claims are generated by InstructGPT Ouyang et al. (2022) and then manually verified. The original benchmark defines two settings: binary classification (supported vs. refuted) and three-class classification (supported, refuted, NEI), but focuses only on label prediction, without providing explanations.
3.2 Our Dataset: Table-Aligned Explanations
Figure 1: Overall dataset creation process.
We extend SciTab by adding cell-level explanations, i.e. table regions required to support or refute each claim. We focus on the supported and refuted claims (868 total), leaving out NEI cases, which are typically under-specified. As shown in Figure 1, our annotation pipeline includes two tasks: claim label verification and evidence selection.
Human Annotation.
Each annotator is given a claim, its label, the associated table, and the caption. Annotators first verify the correctness of the claim. If it is clearly supported or refuted, they mark it as Good; if the claim is unclear, malformed, or unsupported by the table, they choose Do Nothing or optionally revise it (Revised). For Good or Revised claims, annotators highlight the minimal set of table cells required to determine the label. Annotation was performed by four NLP researchers (authors of this paper).
Post-Processing.
We discard samples where all table cells are marked (non-informative) or contain NaN values, resulting in a final dataset of 372 aligned samples (195 supported, 177 refuted).
Inter-Annotator Agreement.
To assess annotation quality, we conducted a second round of labeling on 50 randomly selected tables (covering 137 claims) by a different annotator. Using the first annotation as ground truth, we obtained 75.2% precision, 89.1% recall, and 78.0% macro-F1 for cell-level overlap.
3.3 A Proposed Taxonomy
During annotation, we observed frequent edge cases where claim verification was hindered by poor table formatting, unclear language, or missing context. We propose a taxonomy of five ambiguity types based on annotator notes and discussion:
  • (i) Table Conversion Errors: artifacts introduced during table extraction (e.g. merged cells, missing entries, formatting loss).
  • (ii) Additional Context Requirements: Claims referencing abbreviations, statistical tests, or assumptions not recoverable from the table alone.
  • (iii) Unexpected Claim Types: Descriptive or meta-level claims (e.g., “Table 4 lists the scores of different models.”) that require no reasoning.
  • (iv) Subjective Adjectives: Use of vague or non-quantifiable terms (e.g., “poor”, “substantial”, or “a large margin”).
  • (v) Unclear Claims: Ambiguous references to table elements (e.g., “this model”, “these scores”).
We provide examples for each case in Appendix A. We hope this taxonomy will guide future dataset development and improve robustness in scientific claim verification tasks.
4 Experimental Settings
Models.
We use three groups of models in our experiments: Table-based LLMs, Open-source LLMs, and Closed-source LLMs. For table-based LLMs, we use TAPAS-base and TAPAS-large Herzig et al. (2020), pretrained for reasoning over tabular input. For open-source LLMs, we use the Instruction-tuned variants of Qwen 2.5 (7B and 72B, Yang et al., 2025a) and Llama 3.1 (8B and 70B, Grattafiori et al., 2024). For closed-source LLMs, we use GPT-4o Hurst et al. (2024).
Prompting strategies.
Our dataset contains two subtasks: (1) claim label prediction and (2) cell-level evidence selection. We conduct experiments using three prompting strategies: zero-shot, few-shot, and Chain-of-Thought (CoT; Wei et al., 2022). For few-shot and CoT promptings, we use four demonstration examples selected from the revised subset of samples not included in the evaluation set, ensuring fair evaluation.
Tabular Representation.
Following Wang et al. (2024), who found that the PIPE encoding format with explicit tags (e.g. Col and Row 1) outperformed HTML, TSV, and Markdown formats for tabular data, we adopt PIPE encoding for all of our experiments.
5 Results
Following the original SciTab dataset, we use Macro-F1 for our evaluation. All results are shown in Table 1. It is noted that, due to cost constraints, we only run GPT-4o on 100 samples selected to match the label distribution of the entire dataset.
Model Claim Labeling Cell Selection
Zero Few CoT Zero Few CoT
TAPAS-base 48.1 - - - - -
TAPAS-large 51.6 - - - - -
Llama 3.1 8B 53.2 59.5 62.4 23.6 22.3 22.6
Llama 3.1 70B 75.2 75.0 73.9 31.8 28.8 36.8
Qwen 2.5 7B 66.3 68.1 67.9 20.7 16.6 17.0
Qwen 2.5 72B 83.5 84.7 81.5 32.8 46.7 50.8
GPT-4o 88.4 87.0 88.0 32.4 32.9 34.8
Table 1: Macro-F1 scores of the models on our dataset. ‘Zero’, ‘Few’, and ‘CoT’ denote zero-shot, few-shot, and CoT prompting, respectively.
Claim Prediction Results.
As expected, GPT-4o achieves the highest score. Larger models, such as Qwen 2.5 72B and Llama 3.1 70B, outperform their smaller 7B and 8B counterparts, and all LLMs surpass the performance of the previous table-based model, TAPAS. We also observe that few-shot and CoT prompting are less effective for larger, well-instructed models like the 70B variants and GPT-4o on this familiar label classification task, but remain beneficial for smaller models.
Evidence Selection Results.
Compared to claim label prediction, evidence cell selection is a more challenging task that most LLMs are unfamiliar with. The input consists of a claim and a table, and the output is a list of cell positions—each defined by a row and column index—required to determine the claim’s label. This structured output format adds complexity, and overall, all models struggle to achieve high scores on this task. In the zero-shot setting, GPT-4o, Llama 3.1 70B, and Qwen 2.5 72B achieve comparable scores. Under few-shot and CoT prompting, GPT-4o’s performance remains relatively stable, while Qwen 2.5 72B sees an 18.0 F1 improvement from zero-shot to CoT. CoT prompting also boosts Llama 3.1 70B’s performance. In contrast, smaller models (7B–8B) show decreased performance under both few-shot and CoT prompting compared to zero-shot.
Overall, compared to the human agreement score (78.0 macro F1), the best model still falls short, indicating room for improvement on this task. Despite its difficulty and the possibility of multiple valid reasoning paths, our proposed evidence cells can be seen as a minimal, useful set for claim verification. In the era of black-box LLMs, focusing solely on the final label is insufficient—evidence selection is equally important for explainability. Our work takes a first step toward more interpretable evaluation and highlights the underlying reasoning abilities of models.
Model Table Exp. Table + Exp.
Llama 3.1 8B 53.2 56.9 63.0
Llama 3.1 70B 75.2 80.1 80.9
Qwen 2.5 7B 66.3 67.5 69.8
Qwen 2.5 72B 83.5 80.6 81.9
Table 2: Macro-F1 scores of the models on our dataset using different types of input table contexts. “Exp.” refers to our explanation table cells. For all experiments in this table, we use zero-shot prompting.
Effectiveness of Our Explanation Cells
To evaluate the effectiveness of our explanation cells, we assess models under two different settings: (1) using only our explanation table cells, and (2) using both the original table and our explanation table cells. The results are shown in Table 2. On average, we observe that using only our explanation cells or combining them with the original table leads to improved task performance.
6 Analyses
To better understand the correlation between the claim label prediction task and the cell evidence selection task, we categorize outcomes into four types: Correct–Correct, Correct–Incorrect, Incorrect–Correct, and Incorrect–Incorrect. For claim label prediction, correctness is easily determined based on whether the predicted label (Supported or Refuted) matches the ground truth. In contrast, cell evidence selection involves list-based predictions, making exact matches more challenging. Therefore, we consider two evaluation criteria: exact match (EM) and a relaxed case where an F1 score of 50.0 or higher is considered correct.
Claim Cell L 8B L 70B Q 7B Q 72B GPT
Setting 1: Exact Match for Both Tasks
C C 0.0 0.0 0.0 4.6 0.0
C I 63.4 73.9 68.0 73.4 88.0
I C 0.0 0.0 0.0 0.0 0.0
I I 36.6 26.1 32.0 22.0 12.0
Setting 2: F1 >= 50 in Cell Selection
C C 10.5 26.1 4.3 44.1 30.0
C I 53.0 47.8 63.7 33.9 58.0
I C 5.6 10.5 2.7 8.9 7.0
I I 30.9 15.6 29.3 13.2 5.0
Table 3: Categorical statistics (%) showing the correlation between the claim label prediction and cell evidence selection tasks. C and I denote Correct and Incorrect, respectively. L and Q denote Llama and Qwen, respectively. The results are from CoT prompting.
The percentage distribution of these cases is shown in Table 3. The case where both tasks are correct (C–C) is what we expect. However, as shown in the table, none of the models achieve a percentage of 50% for this case—even in the second setting, where an F1 score of 50.0 or higher is considered correct for the cell selection task. This suggests that while models often predict the claim label correctly, they lack the ability to select the minimal subset of table cells necessary to support that prediction.
7 Conclusion
In this work, we highlighted the limitations of scientific claim verification systems that focus solely on label prediction, arguing for the importance of interpretability through evidence selection. By reframing table-text alignment as an explanation task and introducing a new dataset with human-annotated cell-level rationales, we provide a more rigorous benchmark for evaluating model reasoning. Additionally, we proposed a taxonomy of ambiguous cases in claim verification against tables, which can support future work on dataset construction. Our findings demonstrate that while LLMs often predict the correct claim labels, they frequently fail to identify the minimal supporting evidence, revealing a gap between accuracy and faithful reasoning. This underscores the need for future work to prioritize not just correctness, but also alignment with human-understandable rationales in scientific fact verification tasks.
Limitations
Our work has several limitations. First, the annotation scale is modest, with 868 claims as input and only 372 claims in the final dataset, which may affect the statistical reliability and generalizability of the findings. Second, the dataset originates from a specific domain (computer science), which may limit its applicability to tables and claims from other domains. Third, the PIPE encoding method used may not be well-suited for handling complex table structures, suggesting the need for more robust encoding approaches.
Ethical Statement and Broader Impact
We built our dataset based on the publicly available SciTab dataset, which is released under the MIT License. We respect the terms of this license and provide appropriate attribution to the original authors. To extend the dataset, four NLP researchers manually annotated the data. We created and followed a detailed annotation guideline to ensure consistency, clarity, and fairness in the annotation process. The dataset does not include any personal or sensitive information.
Appendix A Dataset Creation
A.1 A Proposed Taxonomy
We present examples for our proposed taxonomy in Section 3.3 in Tables 4, 5, 6, 7, and 8, respectively.
Claim Comparing POS and SEM tagging (Table 5), we note that higher layer representations do not necessarily improve SEM tagging, while POS tagging does not peak at layer 1. We noticed no improvements in both translation (+0.9 BLEU) and POS and SEM tagging (up to +0.6% accuracy) when using features extracted from an NMT model trained with residual connections (Table 5).
Label Refuted
Table Caption Table 5: POS and SEM tagging accuracy with features from different layers of 4-layer Uni/Bidirectional/Residual NMT encoders, averaged over all non-English target languages.
Table Uni | POS | 0 87.9 | 1 92.0 | 2 91.7 | 3 91.8 | 4 91.9 Uni | SEM | 81.8 | 87.8 | 87.4 | 87.6 | 88.2 Bi | POS | 87.9 | 93.3 | 92.9 | 93.2 | 92.8 Bi | SEM | 81.9 | 91.3 | 90.8 | 91.9 | 91.9 Res | POS | 87.9 | 92.5 | 91.9 | 92.0 | 92.4 Res | SEM | 81.9 | 88.2 | 87.5 | 87.6 | 88.5
Table 4: Example of (i) Table Conversion Errors. The column headers are merged with the data values. For example, "0 87.9" incorrectly combines the column name 0 and the value 87.9.
Claim After removing the graph attention module, our model gives 24.9 BLEU points.
Label Supported
Table Caption Table 9: Ablation study for modules used in the graph encoder and the LSTM decoder
Table [BOLD] Model | B | C DCGCN4 | 25.5 | 55.4 Encoder Modules | [EMPTY] | [EMPTY] -Linear Combination | 23.7 | 53.2 -Global Node | 24.2 | 54.6 -Direction Aggregation | 24.6 | 54.6 -Graph Attention | 24.9 | 54.7 -Global Node &Linear Combination | 22.9 | 52.4 Decoder Modules | [EMPTY] | [EMPTY] -Coverage Mechanism | 23.8 | 53.0
Table 5: Example of (ii) Additional Context Requirements. B and C stand for BLEU and CHRF++, respectively, but this cannot be inferred from the claim, caption, or table alone. It requires additional context from the original paper.
Claim Table 4 lists the EM/F1 score of different models.
Label Supported
Table Caption Table 4: Exact match/F1-score on SQuad dataset. “#Params”: the parameter number of Base. rnet*: results published by Wang et al. (2017).
Table Model | #Params | Base | +Elmo rnet* | - | 71.1/79.5 | -/- LSTM | 2.67M | [BOLD] 70.46/78.98 | 75.17/82.79 GRU | 2.31M | 70.41/ [BOLD] 79.15 | 75.81/83.12 ATR | 1.59M | 69.73/78.70 | 75.06/82.76 SRU | 2.44M | 69.27/78.41 | 74.56/82.50 LRN | 2.14M | 70.11/78.83 | [BOLD] 76.14/ [BOLD] 83.83
Table 6: Example of (iii) Unexpected Claim Types. The claim simply describes what the table shows, similar to the caption, and does not require any reasoning or data to support it.
Claim [CONTINUE] RELIS significantly outperforms the other RL-based systems.
Label Supported
Caption Table 3: Results of non-RL (top), cross-input (DeepTD) and input-specific (REAPER) RL approaches (middle) compared with RELIS.
Table [EMPTY] | DUC’01 R1 | DUC’01 R2 | DUC’02 R1 | DUC’02 R2 | DUC’04 R1 | DUC’04 R2 ICSI | 33.31 | 7.33 | 35.04 | 8.51 | 37.31 | 9.36 PriorSum | 35.98 | 7.89 | 36.63 | 8.97 | 38.91 | 10.07 TCSum | <bold>36.45</bold> | 7.66 | 36.90 | 8.61 | 38.27 | 9.66 TCSum- | 33.45 | 6.07 | 34.02 | 7.39 | 35.66 | 8.66 SRSum | 36.04 | 8.44 | <bold>38.93</bold> | <bold>10.29</bold> | 39.29 | 10.70 DeepTD | 28.74 | 5.95 | 31.63 | 7.09 | 33.57 | 7.96 REAPER | 32.43 | 6.84 | 35.03 | 8.11 | 37.22 | 8.64 RELIS | 34.73 | <bold>8.66</bold> | 37.11 | 9.12 | <bold>39.34</bold> | <bold>10.73</bold>
Table 7: Example of (iv) Subjective Adjectives. Whether the performance is considered “significant” depends on how the term is defined. Moreover, many argue that using the word “significant” requires the result to pass some form of statistical test. The original version of the first row is: [EMPTY] | DUC’01 <italic>R</italic>1 | DUC’01 <italic>R</italic>2 | DUC’02 <italic>R</italic>1 | DUC’02 <italic>R</italic>2 | DUC’04 <italic>R</italic>1 | DUC’04 <italic>R</italic>2.
Claim It closely matches the performance of ORACLE with only 0.40% absolute difference.
Label Supported
Caption Table 3: Accuracy of transferring between aspects. Models with † use labeled data from source aspects. Models with ‡ use human rationales on the target aspect.
Table Source | Target | Svm | Ra-Svm‡ | Ra-Cnn‡ | Trans† | Ra-Trans‡† | Ours‡† | Oracle† Beer aroma+palate | Beer look | 74.41 | 74.83 | 74.94 | 72.75 | 76.41 | [BOLD] 79.53 | 80.29 Beer look+palate | Beer aroma | 68.57 | 69.23 | 67.55 | 69.92 | 76.45 | [BOLD] 77.94 | 78.11 Beer look+aroma | Beer palate | 63.88 | 67.82 | 65.72 | 74.66 | 73.4 | [BOLD] 75.24 | 75.5
Table 8: Example of (v) Unclear Claims. It is unclear what entity the pronoun “it” refers to.